OpenStudy (anonymous):

Which of the following represents the area of a rectangle whose length is x + 1 and whose width is x + 11?

OpenStudy (diyadiya):

Area Of Rectangle =Length x Width =(x+1)(x+11)

OpenStudy (anonymous):

A: x2 + 11x + 12 B: x2 + 10x + 11 C: x2 + 11 D: x2 + 12x + 11

OpenStudy (anonymous):

D?

OpenStudy (diyadiya):

(x+1)(x+11)=x^2+11x+x+11=x^2+12x+11 Yes D

OpenStudy (anonymous):

thanks!